OpenStudy (anonymous):

The average of all odd numbers less than 100 is-----------?

OpenStudy (anonymous):

This is similar to the last deal. You're looking for: [sum of all odd numbers from 1 to 99]/[number of terms] The common difference is 2, and you have 50 terms in the sum, as given by, 99 = 2 + (n-1)*2 -> n = 50 by same reasoning as in your last question. This will be your denominator. Now we need to find the numerator. It will be an arithmetic sum, as before: (n/2)(a[1]+a[50]) = (50/2)*(1+99) = 2500. The average is then, 2500/50 = 50